COMBINATION BANDS OF SMALL CARBON CLUSTERS: AN ISOTOPIC INFRARED STUDY

\underline{JASON F. FULLER}, JAN SZCZEPANSKI and MARTIN VALA , Department of Chemistry and The Center for Chemical Physics, Gainesville, FL 32611.

The ^{12,13}C isotopic substitution method applied to IR absorption spectra has been successful in identifying the size and the geometry of many molecular systems, including carbon clusters.^{a} Using this method we obtained ^{13}C-labeled IR spectra in the ν > 2200 cm^{−1} region for linear carbon clusters isolated in an Ar matrix. By comparison to the calculated isotopomer frequencies at the B3LYP/6-31G* level, the observed bands have been assigned to combination modes of 12,13C_{n} clusters. In addition, from these combination band frequencies the few symmetrical mode vibrations in the linear C_{n} (n< 10) clusters have been deduced.

aA. Van Orden and R. J. Saykally, Chem. Rev. \underline{98}, 2313 (1998), and references therein.